Let A= rolling 2 on pair of fair dice and B = getting 1 heads on 3 rolls of a fair coin. Assuming the rolls of the dice are independent of the coin tosses, what is Prob (A AND B)? (round to 4 decimals)
Answer #1

Ans:

P(A)=(1/6)*(1/6)=1/36

P(B)=3C1*0.5*(1-0.5)^2=0.375

As,A and B are independent.

P(A and B)=P(A) *P(B)(1/36)*0.375=0.0104